4. What is the pH of a 0.350 M solution of lithium citrate? (Ka Citric Acid = 7.4 x 10-4


1
Expert's answer
2022-02-01T23:02:02-0500

Lithium Citrate is salt and this salt go for anionic hydrolysis.

Ka = 7.4 × 10-4 .

Pka = - logka = - log 7.4 × 10-4 = 4 - log 7.4

Pka = 4 - 0.8 = 3.2


So, PH = 7 + Pka /2 + ( logc )/2

So, PH = 7 + 3.2 /2 + ( log 0.350 ) /2

So, pH = 7 + 1.6 + -0.456 /2

So, pH = 8.6 - 0.227 = 8.36